# Topological and nontopological degeneracies in generalized string-net models

### v3

@@ III.2 The output category
By construction, the Drinfeld center $\mathcal{Z(C)}$ of a UFC $\mathcal{C}$ is a UMTC whose objects are built from $\mathcal{C}$. As a UMTC, the output category $\mathcal{Z(C)}$ describes an achiral $(c\mod 8=0)$ anyon theory (see, e.g., Refs. [35, 36, 16, 22] for more informations). As mentioned in Ref. [14], topological orders associated with Drinfeld centers are believed to be the most general class of bosonic topological order compatible with gapped boundaries [40, 11, 41]. One way to build the Drinfeld center of a UFC consists in deriving the Ocneanu tube algebra [31, 32] and finding its irreducible representations. The main lines of this construction are summarized in Appendix A, and some examples are given in Appendix B. In the following, we will denote by Roman letters simple objects of $\mathcal{C}$ and by capital Roman letters simple objects of $\mathcal{Z(C)}$, except for the vacua which are denoted $1$ and ${\bf 1}$, respectively.
@@
Finally, let us mention that an alternative route to the description of the excitations consists in finding operators associated with closed strings that commute with the Hamiltonian (10). Such an approach based on string operators is detailed in Refs. [10, 14], and the relation between the two approaches is discussed in [12].
@@ III.3 Fluxon model
In this work, we will focus on a somewhat simpler Hamiltonian than the full string-net Hamiltonian of Eq. (8), since we shall strictly enforce the vertex constraint. This is essentially equivalent to taking $J_{v}\rightarrow\infty$ in Eq. (8) and considering only low-energy physics. Even more simply, we can think of this constraint as restricting the Hilbert space to include only the states that are +1 eigenstates of $A_{v}$ for all vertices. Within this restricted Hilbert space and setting $J_{p}=1$, the Hamiltonian takes the simpler form
@@
The ground-state space is entirely contained in this restricted Hilbert space. All excitations of the model that remain within this space are plaquette excitations, and known as fluxons. These are the excitations we will focus on in this paper. In Eq. (10), $B_{p}$ is the projector onto the vacuum of $\mathcal{Z(C)}$ in the plaquette $p$. Consequently, the ground-state energy of the system is given by $E_{0}=-N_{\rm p}$, where $N_{\rm p}$ is the total number of plaquettes and a state with $q$ fluxons has an energy $E_{q}=-N_{\rm p}+q=E_{0}+q$.
The Hamiltonian (10) assigns a constant energy penalty of one unit to each plaquette that is not in the vacuum state. It is easy to generalize Eq. (10) to a form that instead assigns a different energy to each different type of fluxon excitation. Similarly, we can also study the case in which the energy penalties depend on which particular plaquette of the lattice we are considering. We briefly address these more general cases in Appendix D.
The Hamiltonian (10) is obviously fine-tuned compared to more generic cases such as that considered in Appendix D. Nonetheless, the techniques we introduce here can be used more broadly and our main point is to establish and test our method.
@@ III.4 Identifying fluxons
We will need to determine which objects from $\cal{Z}(\cal{C})$ are actually fluxons. Using the tube algebra construction, it is straightforward to see that $A$ is a fluxon iff $n_{A,1}\neq 0$ (see Appendix A). This definition also includes the trivial fluxon $A=\mathbf{1}$ (vacuum), although, strictly speaking, it is not an excitation. Other fluxons will be called nontrivial in the following. The set of all fluxons in $\cal{Z}(\cal{C})$ will be denoted $\mathcal{F}$, while that of nontrivial fluxons will be denoted $\mathcal{F}^{*}$. As shown in Appendix C, the vector ${\bf n}_{1}$, the components of which are $n_{A,1}$, obeys
@@
where $S$ and $T$ are the modular matrices of $\mathcal{Z(C)}$. In the following, we will make extensive use of these important relations. Similar equations have appeared in other contexts such as anyon condensation [42] or gapped boundaries [43, 44]. Here they are obtained as properties of fluxons (i.e. plaquette excitations). In addition, if the input category $\mathcal{C}$ is commutative (noncommutative), the number $N_{\mathcal{F}}$ of fluxons equals (is strictly smaller than) the number $N_{\mathcal{C}}$ of simple objects in $\mathcal{C}$. Other simple objects (non fluxons) in $\mathcal{Z(C)}$ are associated with violations of the vertex constraint (see Sec. III.1). At first glance, it may seem strange to worry about these latter excitations since they do not belong to the Hilbert space $\mathcal{H}$. However, we will see that they do play a role in the degeneracies of the energy spectrum.
Finally, let us comment briefly on Morita-equivalent categories. As mentioned in the Introduction, two different input categories may have the same Drinfeld center. Thus, one could naively think that two string-net models built from these two categories have the same plaquette excitations but, in general, this is not true. Indeed, as discussed above, multiplicities $n_{A,s}$ do not depend only on $\mathcal{Z}(\mathcal{C})$, but also on $\mathcal{C}$. We will illustrate such a situation in Sec. VI.1.
@@ IV Spectral degeneracies
The goal of the present paper is to explain how to compute the degeneracy of the $q$-fluxon energy level in a string-net model. As can be anticipated, this degeneracy depends on the surface topology and on the input UFC, which, as discussed above, determines the nature of the excitations. Before proceeding further, let us stress that this degeneracy has recently been computed in Ref. [25] for any modular input UFC. Here, we aim at giving a general expression valid for any input UFC (see Fig. 4 for an overview).
@@ IV.1 Moore-Seiberg-Banks formula
Another special limit of Eq. (13) is the case in which there are three punctures in a $g=0$ spherical surface (see again Fig. 7). In this case, one recovers precisely the Verlinde formula (6). Note that Eq. (13) has a useful property that if one of the punctures $j$ is labeled by the vaccum, i.e., if $A_{j}=\mathbf{1}$, it can simply be dropped from the list. So, for example, one has
@@ IV.2 Application to string-net fluxon model
From the previous Sec. IV.1, one might guess that the corresponding degeneracy is given by $\dim_{{\cal Z}({\cal C})}(g;A_{1},\ldots,A_{q})$ defined in Eq. (13), where the UMTC in question here is ${\cal Z}({\cal C})$. In the case in which $\cal C$ has commutative fusion rules, this guess is in fact correct. However, more generally, the number of states built from the category $\cal C$ on a surface of genus $g$ with fluxons $A_{i}$ through plaquettes $i=1\ldots q$ is given by
@@
where we used Eqs. (16) and (13) to go from the first to the second line, and Eq. (11) to go from the second to the third. By adding the proper local operator to the Hamiltonian, it is possible to identify the fluxon labels present in the different plaquettes. Therefore, this total degeneracy carries some additional nontopological degeneracy with respect to Eq. (16). Thus in the presence of small amounts of “disorder,” the degeneracy calculated in Eq. (17) is genericallly lifted. The above equation for the degeneracy of a level with $q$-fluxons and involving the internal multiplicities $n_{A,1}$ is the main result of the present work.
@@ IV.2.1 Manifold with boundaries
We now extend the previous results to string-net models on surfaces with boundaries. For a given topological order, there are several types of possible gapped boundaries [40, 44]. Here, we only consider smooth boundaries corresponding to the condensation of the fluxons. Consider starting with an orientable manifold without boundary of genus $g$. Then we can imagine poking $b$ (potentially large) holes in this manifold to obtain a manifold with boundaries. These holes are essentially (potentially large) plaquettes themselves — the only distinction between a boundary hole and a regular plaquette is that in Eq. (10) we sum over only plaquettes that are not these boundary holes, i.e., the energy cost of a boundary hole is zero.
@@
Note that, apart from the substitution of $b$ with $q$, the only difference between Eq. (18) and Eq. (17) comes from the fact that we sum either on $\mathcal{F}$ or on $\mathcal{F^{*}}$. Equation (18) should be considered as a generalization of Eq. (14) to the case with boundaries and with internal multiplicities. In the absence of boundaries, i.e., for $b=0$, one recovers Eq. (14):
@@ IV.2.2 Total Hilbert-space dimension
One sees straightforwardly that for commutative UFCs ($n_{A,1}=0$ or $1$), this dimension only depends on the number of vertices, whereas in the noncommutative case for which $n_{A,1}$ can be larger than 1, $\dim\mathcal{H}$ also depends on $N_{\text{p}}$. Hence, in this latter case, the Hilbert-space dimension is sensitive to the surface topology. If the input category $\mathcal{C}$ is a UMTC, one recovers the result of [46, 25]:
@@
where $\mathcal{D}=\mathcal{D}_{\mathcal{C}}^{2}$ is the total quantum dimension of $\mathcal{Z(C)}$, and $\mathcal{D}_{\mathcal{C}}$ is that of $\mathcal{C}$. Replacing $N_{\text{p}}+b$ by $22-2g+N_{\text{v}}/$ 2 and keeping in mind that $S_{\mathbf{1},A}=d_{A}/\mathcal{D}$ and $n_{A,1}\leq d_{A}$, one gets, in the thermodynamic limit,
@@ V.1 Modular categories
When the input category of the string-net model is modular, it is possible to compute the degeneracies of the excited levels by using only the fusion properties of the input category. The derivation of this equation is presented in Ref. [25], where the partition function is computed from a microscopic point of view for string-net models built from an input category $\mathcal{C}$ which is a UMTC. We aim here at making the link between the results of the previous sections and the ones of Ref. [25]. In the case in which the input category $\mathcal{C}$ is a UMTC, the Drinfeld center is the product of two copies of $\mathcal{C}$ of opposite chiralities [47]:
@@
In this case, any object $A$ of the center can be represented by a couple $(i,j)$ of objects $i\in\mathcal{C}$ and $j\in\bar{\mathcal{C}}$. The $S$-matrix of the Drinfeld center $\mathcal{Z}(\mathcal{C})$ has a simple expression in terms of the $S$-matrix of the input category $\mathcal{C}$ (for clarity, we will denote them as, respectively, $S$ and $s$):
@@ V.2 Commutative categories
The previous equations do not hold in the case in which the input category is commutative but not modular, since we cannot define a modular $S$-matrix. However, for any commutative input category, we can always find a unitary matrix $\tilde{s}$, called the mock $S$-matrix, that simultaneously diagonalizes its fusion matrices $N_{i}$. A Verlinde-like formula (6) holds between the $N_{i}$ and $\tilde{s}$ [see Eq. (87)]. The columns of the matrix $\tilde{s}$ can be indexed by the fluxons $A$ in $\mathcal{Z}(\mathcal{C})$ (for details on the mock $S$-matrix and the “mixed” notation for the matrix elements $\tilde{s}_{i,A}$ with $i\in\mathcal{C}$ and $A\in\mathcal{F}$, see Appendix C.1.1). In particular, if $A$ is a fluxon, $S_{\mathbf{1},A}=d_{A}/\mathcal{D}=\tilde{s}_{1,A}^{2}$. Also, for any commutative UFC, the multiplicities $n_{A,1}$ only take values 0 or 1, and the number of fluxons is equal to the number of simple objects in the input category. With these insights at hand, we can rewrite Eq. (23) as
@@
Generally speaking, it is not possible to compute the ground-state degeneracy $D_{{\cal C}}(g,0,0)$ simply from the knowledge of the fusion rules of the input category. In fact, two input categories that have the same fusion rules but different $F$-symbols generally lead to two different Drinfeld centers. In particular, the number of simple objects of the Drinfeld centers need not be the same [48]. And even if the numbers match, the corresponding quantum dimensions $d_{A}$ need not be equal, and therefore the level degeneracies can differ. However, there is still a way to write $D_{{\cal C}}(g,0,0)$ only in terms of the input data. This way, which is a bit cumbersome, requires the use of the $F$-symbols of the category (see Ref. [29]).
@@ VI Examples
This section is devoted to non-trivial examples that illustrate the formulas obtained in the previous sections. Before discussing some specific examples, let us give some general results that are valid for any input category ${\cal C}$. In Table 1, we give the degeneracies $D_{{\cal C}}(g,b,q)$ [see Eq. (23)] of the first three energy levels for various simple surface topologies.
@@ VI.1 Rep($S_{3}$) and Vec($S_{3}$) categories
We present here two string-net models built from two Morita-equivalent categories: Rep($S_{3}$) and Vec($S_{3}$), where $S_{3}$ is the symmetric group on a set of three elements (symmetry group of the oriented equilateral triangle). They correspond, respectively, to the category of irreducible representations of $S_{3}$ and the category of the elements of the group $S_{3}$.
The category $\mathcal{C}$ = Rep($S_{3}$) contains $N_{\mathcal{C}}=3$ simple objects $\{1,2,3\}$ with quantum dimensions $\{1,1,2\}$, which correspond to the irreducible representations of $S_{3}$. Fusion rules and $F$-symbols can be found, e.g., in Ref. [10]. The fusion rules are commutative, but the category is non-Abelian, braided, and non modular.
The category $\mathcal{C}$ = Vec($S_{3}$) contains $N_{\mathcal{C}}=6$ simple objects $\{e,\zeta,\zeta^{2},\tau,\tau\zeta,\tau\zeta^{2}\}$, which are the elements of the group $S_{3}$, ($e$ is the identity element, $\zeta$ is a $32\pi/$ 3 rotation, and $\tau$ is a mirror). The fusion rules are simply the multiplication rules of the group. All elements have quantum dimension $d_{s}=1$. The fusion rules are noncommutative, but the category is Abelian.
The two categories are Morita-equivalent and lead to the same Drinfeld center $\mathcal{Z}(S_{3})$. The latter contains $N_{\mathcal{Z}}=8$ simple objects, denoted by $\{A,B,C,D,E,F,G,H\}$; see Ref. [49]. The quantum dimensions are $\{1,1,2,3,3,2,2,2\}$, and the total quantum dimension is $\mathcal{D}=6$.
While the two categories share the same center $\mathcal{Z}(S_{3})$, the tube algebras are different (see Appendix B). In particular, the internal multiplicities of the quasiparticles, $n_{J,s}$, differ. We give them as rectangular matrices $n_{\mathcal{C}}$ with $N_{\mathcal{C}}$ columns and $N_{\mathcal{Z}}$ rows:
@@
[…]

### v4

@@ III.2 The output category
By construction, the Drinfeld center $\mathcal{Z(C)}$ of a UFC $\mathcal{C}$ is a UMTC whose objects are built from $\mathcal{C}$. As a UMTC, the output category $\mathcal{Z(C)}$ describes an achiral $(c\mod 8=0)$ anyon theory (see, e.g., Refs. [35, 36, 16, 22] for more informations). As mentioned in Ref. [14], topological order associated with Drinfeld centers are believed to be the most general class of bosonic topological order compatible with gapped boundaries [40, 11, 41]. One way to build the Drinfeld center of a UFC consists in deriving the Ocneanu tube algebra [31, 32] and finding its irreducible representations. The main lines of this construction are summarized in Appendix A and some examples are given in Appendix B. In the following, we will denote by Roman letters simple objects of $\mathcal{C}$ and by capital Roman letters simple objects of $\mathcal{Z(C)}$, except for the vacua which are denoted $1$ and ${\bf 1}$, respectively.
@@
Finally, let us mention that an alternative route to the description of the excitations consists in finding operators associated to closed strings which commute with the Hamiltonian (10). Such an approach based on string operators is detailed in Refs. [10, 14] and the relation between the two approaches is discussed in [12].
@@ III.3 Fluxon model
In this work, we will focus on a somewhat simpler Hamiltonian than the full string-net Hamiltonian of Eq. (8), since we shall strictly enforce the vertex constraint. This is essentially equivalent to taking $J_{v}\rightarrow\infty$ in Eq. (8) and considering only low-energy physics. Even more simply we can think of this constraint as restricting the Hilbert space to include only the states that are +1 eigenstates of $A_{v}$ for all vertices. Within this restricted Hilbert space and setting $J_{p}=1$, the Hamiltonian takes the simpler form
@@
The ground-state space is entirely contained in this restricted Hilbert space. All excitations of the model that remain within this space are plaquette excitations, and known as fluxons. These are the excitations we will focus on in this paper. In Eq. (10), $B_{p}$ is the projector onto the vacuum of $\mathcal{Z(C)}$ in the plaquette $p$. Consequently, the ground-state energy of the system is given by $E_{0}=-N_{\rm p}$ where $N_{\rm p}$ is the total number of plaquette and a state with $q$ fluxons has an energy $E_{q}=-N_{\rm p}+q=E_{0}+q$.
The Hamiltonian (10) assigns a constant energy penalty of one unit to each plaquette which is not in the vacuum state. It is easy to generalize Eq. (10) to a form that instead assigns a different energy to each different type of fluxon excitation. Similarly, we can also study the case where the energy penalties depend on which particular plaquette of the lattice we are considering. We briefly address these more general cases in Appendix D.
The Hamiltonian (10) is obviously fine tuned compared to more generic cases such as that considered in Appendix D. Nonetheless, the techniques we introduce here can be used more widely and our main point is to establish and test our method.
@@ III.4 Identifying fluxons
We will need to determine which objects from $\cal{Z}(\cal{C})$ are actually fluxons. Using the tube algebra construction, it is straightforward to see that $A$ is a fluxon iff $n_{A,1}\neq 0$ (see Appendix A). This definition also includes the trivial fluxon $A=\mathbf{1}$ (vacuum) although, strictly speaking, it is not an excitation. Other fluxons will be called nontrivial in the following. The set of all fluxons in $\cal{Z}(\cal{C})$ will be denoted $\mathcal{F}$, while that of nontrivial fluxons will be denoted $\mathcal{F}^{*}$. As shown in Appendix C, the vector ${\bf n}_{1}$, the components of which are $n_{A,1}$, obeys
@@
where $S$ and $T$ are the modular matrices of $\mathcal{Z(C)}$. In the following, we will make an extensive use of these important relations. Similar equations have appeared in other contexts such as anyon condensation [42] or gapped boundaries [43, 44]. Here they are obtained as properties of fluxons (i.e. plaquette excitations). In addition, if the input category $\mathcal{C}$ is commutative (noncommutative), the number $N_{\mathcal{F}}$ of fluxons equals (is strictly smaller than) the number $N_{\mathcal{C}}$ of simple objects in $\mathcal{C}$. Other simple objects (non fluxons) in $\mathcal{Z(C)}$ are associated to violations of the vertex constraint (see Sec. III.1). At first glance, it may seem strange to worry about these latter excitations since they do not belong to the Hilbert space $\mathcal{H}$. However, we will see that they do play a role in the degeneracies of the energy spectrum.
Finally, let us make a short comment on Morita-equivalent categories. As mentioned in the introduction, two different input categories may have the same Drinfeld center. Thus, one could naively think that two string-net models built from these two categories have the same plaquette excitations but, in general, this is not true. Indeed, as discussed above, multiplicities $n_{A,s}$ do not depend only on $\mathcal{Z}(\mathcal{C})$, but also on $\mathcal{C}$. We will illustrate such a situation in Sec. VI.1.
@@ IV Spectral degeneracies
The goal of the present paper is to explain how to compute the degeneracy of the $q$-fluxon energy level in a string-net model. As can be anticipated, this degeneracy depends on the surface topology and on the input UFC which, as discussed above, determines the nature of the excitations. Before proceeding further, let us stress that this degeneracy has recently been computed in Ref. [25] for any modular input UFC. Here, we aim at giving a general expression valid for any input UFC (see Fig. 4 for an overview).
@@ IV.1 Moore-Seiberg-Banks formula
Another special limit of Eq. (13) is the case where there are three punctures in a $g=0$ spherical surface (see again Fig. 7). In this case, one recovers precisely the Verlinde formula (6). Note that Eq. (13) has a useful property that if one of the punctures $j$ is labeled by the vaccum, i.e., if $A_{j}=\mathbf{1}$, it can simply be dropped from the list. So, for example, one has
@@ IV.2 Application to string-net fluxon model
From the previous Sec. IV.1, one might guess that the corresponding degeneracy is given by $\dim_{{\cal Z}({\cal C})}(g;A_{1},\ldots,A_{q})$ defined in Eq. (13), where the UMTC in question here is ${\cal Z}({\cal C})$. In the case where $\cal C$ has commutative fusion rules, this guess is in fact correct. However, more generally, the number of states built from the category $\cal C$ on a surface of genus $g$ with fluxons $A_{i}$ through plaquettes $i=1\ldots q$ is given by
@@
where we used Eqs. (16) and (13) to go from the first to the second line and Eq. (11) from the second to the third. By adding the proper local operator to the Hamiltonian, it is possible to identify the fluxon labels present in the different plaquettes. Therefore, this total degeneracy carries some additional nontopological degeneracy with respect to Eq. (16). Thus in the presence of small amounts of “disorder”, the degeneracy calculated in Eq. (17) is genericallly lifted. The above equation for the degeneracy of a level with $q$-fluxons and involving the internal multiplicities $n_{A,1}$ is the main result of the present work.
@@ IV.2.1 Manifold with boundaries
We now extend the previous results to string-net models on surfaces with boundaries. For a given topological order, there are several types of possible gapped boundaries [40, 44]. Here, we only consider smooth boundaries corresponding to the condensation of the fluxons. Consider starting with an orientable manifold without boundary of genus $g$. Then we can imagine poking $b$ (potentially large) holes in this manifold to obtain a manifold with boundary. These holes are essentially (potentially large) plaquettes themselves — the only distinction between a boundary hole and a regular plaquette is that in Eq. (10) we sum over only plaquettes that are not these boundary holes, i.e., the energy cost of a boundary hole is zero.
@@
Note that, apart from the substitution of $b$ with $q$, the only difference between Eq. (18) and Eq. (17) comes from the fact that we sum either on $\mathcal{F}$ or on $\mathcal{F^{*}}$. Equation (18) should be considered as a generalization of Eq. (14) to the case with boundaries and with internal multiplicities. In the absence of boundary, i.e., for $b=0$, one recovers Eq. (14):
@@ IV.2.2 Total Hilbert-space dimension
One straightforwardly sees that for commutative UFCs ($n_{A,1}=0$ or $1$), this dimension only depends on the number of vertices whereas, in the noncommutative case for which $n_{A,1}$ can be larger than 1, $\dim\mathcal{H}$ also depends on $N_{\text{p}}$. Hence, in this latter case, the Hilbert-space dimension is sensitive to the surface topology. If the input category $\mathcal{C}$ is a UMTC, one recovers the result of [46, 25]:
@@
where $\mathcal{D}=\mathcal{D}_{\mathcal{C}}^{2}$ is the total quantum dimension of $\mathcal{Z(C)}$ and $\mathcal{D}_{\mathcal{C}}$ that of $\mathcal{C}$. Replacing $N_{\text{p}}+b$ by $22-2g+N_{\text{v}}/$ 2 and keeping in mind that $S_{\mathbf{1},A}=d_{A}/\mathcal{D}$ and $n_{A,1}\leq d_{A}$, one gets, in the thermodynamic limit:
@@ V.1 Modular categories
When the input category of the string-net model is modular, it is possible to compute the degeneracies of the excited levels by using only the fusion properties of the input category. The derivation of this equation is presented in Ref. [25], where the partition function is computed from a microscopic point of view for string-net models built from an input category $\mathcal{C}$ which is a UMTC. We aim here at making the link between the results of the previous sections and the ones of Ref. [25]. In the case where the input category $\mathcal{C}$ is a UMTC, the Drinfeld center is the product of two copies of $\mathcal{C}$ of opposite chiralities [47]:
@@
In this case, any object $A$ of the center can be represented by a couple $(i,j)$ of objects $i\in\mathcal{C}$ and $j\in\bar{\mathcal{C}}$. The $S$-matrix of the Drinfeld center $\mathcal{Z}(\mathcal{C})$ has a simple expression in terms of the $S$-matrix of the input category $\mathcal{C}$ (for clarity, we will denote them respectively $S$ and $s$):
@@ V.2 Commutative categories
The previous equations do not hold in the case where the input category is commutative but not modular, since we cannot define a modular $S$-matrix. However, for any commutative input category, we can always find a unitary matrix $\tilde{s}$, called the mock $S$-matrix, that simultaneously diagonalizes its fusion matrices $N_{i}$. A Verlinde-like formula (6) holds between the $N_{i}$ and $\tilde{s}$, see Eq. (87). The columns of the matrix $\tilde{s}$ can be indexed by the fluxons $A$ in $\mathcal{Z}(\mathcal{C})$ (for details on the mock $S$-matrix and the “mixed” notation for the matrix elements $\tilde{s}_{i,A}$ with $i\in\mathcal{C}$ and $A\in\mathcal{F}$, see Appendix C.1.1). In particular, if $A$ is a fluxon, $S_{\mathbf{1},A}=d_{A}/\mathcal{D}=\tilde{s}_{1,A}^{2}$. Also, for any commutative UFC, the multiplicities $n_{A,1}$ only take values 0 or 1, and the number of fluxons is equal to the number of simple objects in the input category. With these insights at hand, we can rewrite Eq. (23) as
@@
Generally speaking, it is not possible to compute the ground-state degeneracy $D_{{\cal C}}(g,0,0)$ simply from the knowledge of the fusion rules of the input category. In fact, two input categories that have the same fusion rules but different $F$-symbols generally lead to two different Drinfeld centers. In particular, the number of simple objects of the Drinfeld centers need not be the same [48]. And even if the numbers match, the corresponding quantum dimensions $d_{A}$ need not be equal and therefore the level degeneracies can differ. However, there still is a way to write $D_{{\cal C}}(g,0,0)$ only in terms of the input data. This way, which is a bit cumbersome, requires the use of the $F$-symbols of the category (see Ref. [29]).
@@ VI Examples
This section is devoted to nontrivial examples that illustrate the formulas obtained in the previous sections. Before discussing some specific examples, let us give some general results that are valid for any input category ${\cal C}$. In Table 1, we give the degeneracies $D_{{\cal C}}(g,b,q)$ [see Eq. (23)] of the first three energy levels for various simple surface topologies.
@@ VI.1 Rep($S_{3}$) and Vec($S_{3}$) categories
We present here two string-net models built from two Morita-equivalent categories: Rep($S_{3}$) and Vec($S_{3}$), where $S_{3}$ is the symmetric group on a set of 3 elements (symmetry group of the equilateral triangle). They correspond respectively to the category of irreducible representations of $S_{3}$ and the category of the elements of the group $S_{3}$.
The category $\mathcal{C}$ = Rep($S_{3}$) contains $N_{\mathcal{C}}=3$ simple objects $\{1,2,3\}$ with quantum dimensions $\{1,1,2\}$, which correspond to the irreducible representations of $S_{3}$. Fusion rules and $F$-symbols can be found, e.g., in Ref. [10]. The fusion rules are commutative but the category is non-Abelian, braided, and non modular.
The category $\mathcal{C}$ = Vec($S_{3}$) contains $N_{\mathcal{C}}=6$ simple objects $\{e,\zeta,\zeta^{2},\tau,\tau\zeta,\tau\zeta^{2}\}$, which are the elements of the group $S_{3}$, ($e$ is the identity element, $\zeta$ is a $32\pi/$ 3 rotation and $\tau$ is a mirror). The fusion rules are simply the multiplication rules of the group. All elements have quantum dimension $d_{s}=1$. The fusion rules are noncommutative but the category is Abelian.
The two categories are Morita-equivalent and lead to the same Drinfeld center $\mathcal{Z}(S_{3})$. The latter contains $N_{\mathcal{Z}}=8$ simple objects, denoted by $\{A,B,C,D,E,F,G,H\}$, see Ref. [49]. The quantum dimensions are $\{1,1,2,3,3,2,2,2\}$, and the total quantum dimension is $\mathcal{D}=6$.
While the two categories share the same center $\mathcal{Z}(S_{3})$, the tube algebras are different (see Appendix B). In particular, the internal multiplicities of the quasiparticles, $n_{J,s}$ differ. We give them as rectangular matrices $n_{\mathcal{C}}$ with $N_{\mathcal{C}}$ columns and $N_{\mathcal{Z}}$ rows:
@@
VI.2 Tambara-Yamagami category TY3
The Tambara-Yamagami category for $\mathds{Z}_{3}$ (notated TY3) is interesting as it breaks the tetrahedral symmetry and therefore can only be realized by the generalized string-net model (see for example [14]) and not by the original Levin-Wen model [10]. It is commutative (but non-Abelian and nonbraided), so that $n_{A,1}=1$ for all fluxons. It has $N_{\mathcal{C}}=4$ simple objects $\{1,2,3,\sigma\}$, with quantum dimensions $\{1,1,1,\sqrt{3}\}$. There are two different solutions to the pentagonal equations indexed by $p=\pm 1$ (see [14] for more details on the $F$-symbols). The Drinfeld center $\mathcal{Z}({\rm TY}_{3})$ contains $N_{\mathcal{Z}}=15$ elements $\{1,...,15\}$ with dimensions $\{1,1,1,1,1,1,2,2,2,\sqrt{3},\sqrt{3},\sqrt{3},\sqrt{3},\sqrt{3},\sqrt{3}\}$ and the total quantum dimension is $\mathcal{D}=6$. For the $p=1$ model, the internal multiplicities are: